OpenStudy (anonymous):

Two trains leave stations 238 miles apart at the same time and travel toward each other. One train travels at 90 miles per hour while the other travels at 80 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

1.4 hours to meet.

OpenStudy (anonymous):

thank-you! can you show me how you worked that out please

OpenStudy (anonymous):

When the trains travel towards each other, we add the speeds. Therefore, using the speed- distance formula : Speed = Distance/ Time. 90+80 = 238/ time Time= t t= 238/ 170 Which is t = 7/ 5 which is 1.4 hours. =]

OpenStudy (anonymous):

thank-you soo much

OpenStudy (anonymous):

You're welcome. =D Have a great day.